OpenStudy (anonymous):

if 4 times a number is decreased by 14 the result is 18 find the number

OpenStudy (anonymous):

number=8

OpenStudy (anonymous):

4(x-14)=18 x=32

OpenStudy (anonymous):

4x-14=8 4x=22 x=11/2